OpenStudy (anonymous):

Find the solution to the differential equation dx/dt=2x+2t+tx+4 that satisfies the conditions x>=0 and x(1)=2

OpenStudy (amistre64):

i cant even tell if im reading it right

OpenStudy (amistre64):

combine all your "x" terms

OpenStudy (amistre64):

they will tell you the function associated with "x"

OpenStudy (jamesj):

Notice that 2x+2t+tx+4 = (x+2)(t+2)

OpenStudy (amistre64):

x' =(2+t)x+(2t+4) x'-(2+t)x =2t+4 might be a way to format it to a usual look

OpenStudy (amistre64):

dunno if its best, but at least it gets you to a "pattern" that normal

OpenStudy (jamesj):

Look ... because dx/dt = 2x+2t+tx+4 = (x+2)(t+2) this equation is separable: \[ \frac{dx}{x+2} = (t+2) dt \] Now integrate

OpenStudy (amistre64):

separable is soooo last chapter tho :)

OpenStudy (amistre64):

that was weird

OpenStudy (jamesj):

@Lammy, making sense?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

Ln(x+2)+C = 1/2t^2+2t+c

OpenStudy (anonymous):

C=2ln2-(5/2) so ln(x+2)=1/2t^2+2t+2ln(2)-(5/2)

OpenStudy (amistre64):

only one side needs a +c ; as a shortcut; then solve for x

OpenStudy (amistre64):

ln(x+2) = 1/2 t^2 + 2t + c x+2 = exp(1/2 t^2 + 2t + c) x = exp(1/2 t^2 + 2t + c) - 2 x = Cexp(1/2 t^2 + 2t) - 2

OpenStudy (anonymous):

i dont have to solve for C first?

OpenStudy (jamesj):

You can solve for C now using the initial condition provided to you in the problem: x(1) = 2. So take the general solution \[ x(t) = Ce^{\frac{t^2}{2} + 2t} - 2 \] and now find C.